OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 0 −3 3 2 answer :

OpenStudy (anonymous):

it doesnt have a solution

OpenStudy (whpalmer4):

You need to solve the system of equations. I suggest you do this one by substitution, as it is easy to solve the second equation for \(x\). Rearrange the second equation so \(x\) is by itself on the left. Then rewrite the first equation, substituting whatever is on the right hand side of the equals sign in the second equation after you've rearranged it wherever you find \(x\) in the first equation. Solve that equation for the value of \(y\), then plug the value you get for \(y\) back into the substitution to find \(x\).

OpenStudy (whpalmer4):

@TjRoDz Now what makes you say that?

OpenStudy (whpalmer4):

Here's an example of substitution, but the values will NOT be the same as in this problem, this is only for illustration. \[2x+2y = 6\]\[x-y=3\] Rearrange the second equation to get \(x\) all by itself on the left hand side. Do this by adding \(y\) to both sides: \[x-y+y=3+y\]\[x=3+y\]Now rewrite the first equation, except wherever you see \(x\) you are going to write \((3+y)\): \[2(3+y) + 2y = 6\]Distribute the left hand side \[2*3 + 2*y + 2y = 6\]\[6 + 4y = 6\]Solve for \(y\) by subtracting 6 from both sides \[6-6+4y = 6-6\]\[4y=0\]Solve for \(y\) by dividing both sides by 4\]\[4y/4 = 0/4\]\[y=0\]Now find \(x\) \[x=3+y\]\[x=3+0\]\[x=3\]

OpenStudy (whpalmer4):

Remember, this is a different problem that I've done, don't assume that the answer is the same as for yours!

OpenStudy (anonymous):

ya i dont like math as much so its kind of hard for me so ill try

OpenStudy (anonymous):

y=0 x=3